OpenStudy (anonymous):

(2x+19)-9x=9(13-x)+21 What number does x represent? In other words solve for x

OpenStudy (anonymous):

My bad its 2(2x+19)-9x=9(13-x)=21

OpenStudy (anonymous):

i changed the problem.^^^

OpenStudy (tester97):

hold on

OpenStudy (zubhanwc3):

is it + 9(13-x) or +

OpenStudy (zubhanwc3):

cause if its +, x=67/7

OpenStudy (tester97):

Yea because the equal signs doesnt make sense in the equation

OpenStudy (anonymous):

I just got the answer. Wanna see my work? Just to check my work?